how long can a 100 watt bulb burn with 1kg of any material??"
1
Expert's answer
2017-02-21T08:19:07-0500
Einstein formula is

E=mc^2

Amount of energy in 1kg is

E=1*(3*10^8)^2 = 9*10^16 J

Hence it can burn

9*10^16/100 = 9*10^14 seconds. As one year is 31*10^6 seconds, it will approximately burn 29*10^6 years